OpenStudy (anonymous):

The radius of one sphere is twice as great as the radius of a second sphere. The ratio of the volume of the smaller sphere to the volume of the larger sphere is _____. A.) 1:2 B.) 1:4 C.) 1:8 D.) 1:16

OpenStudy (anonymous):

I don't have a good idea on how to do this, help would be much appreciated! thank you :)

OpenStudy (anonymous):

Do you know the formula for volume of sphere?

OpenStudy (anonymous):

yeah, V=4/3pie r^3

OpenStudy (anonymous):

So R = 2r 4/3pie r^3/4/3pie (2r)^3 = 1/8

OpenStudy (anonymous):

r^3/(2r)^3= 1/2^3 = 1/8

OpenStudy (anonymous):

oh wait, I'm getting confused, so I just plug in 2r for the radius, cube it, then times it by 4/3?

OpenStudy (anonymous):